Calculate the pH of 0.01 M aqueous solution of sodium formate at 298 K. [Given:




Ka




(HCOOH) =1.7 × 10−4 at 298K]

1
Expert's answer
2022-04-04T17:10:05-0400

Solution:

HCOONa(aq) + H2O(l) ⇌ NaOH(aq) + HCOOH(aq)

HCOO(aq) + H2O(l) ⇌ OH(aq) + HCOOH(aq)


ICE Table:



According to the ICE Table:

[OH] = [HCOOH] = 0.01h

[HCOO] = 0.01 × (1 − h)


The equlibrium-constant (K) expression for the reaction is:



Therefore,

Ka × [HCOOH] × [OH] = Kw × [HCOO]

Ka × (0.01h) × (0.01h) = Kw × [0.01 × (1 − h)]

Assume that h << 1, so 0.01 × (1 − h) ≈ 0.01

Thus,

Ka × (0.012h2) = Kw × (0.01 )

(1.7×10−4) × (0.012h2) = (1.0×10−14) × (0.01 )

h2 = 5.88×10−9

h = (5.88×10−9)1/2 = 7.67×10−5

h = 7.67×10−5


[OH] = 0.01h = (0.01) × (7.67×10−5) = 7.67×10−7

pOH = −log[OH] = −log(7.67×10−7) = 6.12

pH = 14 − pOH = 14 − 6.12 = 7.88

pH = 7.88


Answer: pH = 7.88